OpenStudy (anonymous):

Solve the equation with any method -2x^2+x+15=0

OpenStudy (anonymous):

The equation can be written as 2x^2 - x - 15 = 0 Factor, you get (2x + 5)(x + 3) = 0 x = -5/2 or x = -3